Use the given information to find the

​p-value.

​Also, use a 0.05 significance level and state the conclusion about the null hypothesis​ (reject the null hypothesis or fail to reject the null​ hypothesis).

With

Upper H 1

​:

pgreater than

​0.554,

the test statistic is

zequals

1.34.

A.

​0.0901; reject the null hypothesis

B.

​0.0901; fail to reject the null hypothesis

C.

​0.9099; fail to reject the null hypothesis

D.

​0.1802; reject the null hypothesis
Answer #1

SOLUTION:

From given data,

Use the given information to find the ​p-value. ​Also, use a 0.05 significance level and state the conclusion about the null hypothesis​ (reject the null hypothesis or fail to reject the null​ hypothesis).With Upper H 1 : p greater than ​0.554, the test statistic is  z equals 1.34.

Where,

significance level = = 0.05

test statistic = z =1.34

Test hypothesis:

H0 : p = 0.554 (null hypothesis)

H1 : p > 0.554 (Alternative hypothesis)

Using standard normal table,

P-value:

P-value = P(Z > 1.34)

P-value = 1 - P(Z < 1.34)

P-value = 1 - 0.90988

P-value = 0.0901

P-value = 0.0901 > significance level = = 0.05 then

fail to reject the null hypothesis.

Answer : Option (B)

​0.0901; fail to reject the null hypothesis

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.